# On Probabilistic Pullback Metrics on Latent Hyperbolic Manifolds

### v2

@@ 3 Metrics of Hyperbolic LVMs
Tosi et al. (2014) introduced the pullback metric for Euclidean GPLVMs and its use to compute geodesics that adhere to the data distribution. Here, we extend this approach to hyperbolic LVMs, specifically to the GPHLVM introduced in Sec. 2. We show that the resulting metric can be computed similarly to the Euclidean case, with two key exceptions: (1) the pullback metric must be adequately projected onto appropriate tangent spaces; and (2) the computation of the derivatives for the hyperbolic SE kernels is significantly more challenging than in the Euclidean scenario.
@@ 3.1 A General Hyperbolic Pullback Metric
A deterministic immersion ${f:\mathbb{R}^{D_{x}}\to\mathbb{R}^{D_{y}}}$ pulls back the metric into the latent space following (7), i.e., ${\bm{G}^{\text{P},\mathbb{R}}_{\bm{x}}=\bm{J}^{\mathsf{T}}\bm{J}}$. In the context of LVMs, the immersion $f$ is stochastic. As developed in (Tosi et al., 2014), the stochastic immersion $f$ induces a distribution over its Jacobian $\bm{J}$, which itself induces a distribution over the pullback metric. As Tosi et al. (2014), we consider LVMs for which the probability over $\bm{J}$ follows a Gaussian distribution. Assuming independent rows $\bm{J}_{d}\in\mathbb{R}^{D_{x}}$, each with its own mean but shared covariance matrix, the Jacobian distribution is of the form,
$p(\bm{J})=\prod_{d=1}^{D_{y}}\mathcal{N}(\bm{J}_{d}\mid\bm{\mu}_{d},\bm{\Sigma}_{J})\,.$ (8)
Therefore, the metric tensor $\bm{G}^{\text{P},\mathbb{R}}_{\bm{x}}$ follows a non-central Wishart distribution (Anderson, 1946),
@@
While the hyperbolic case follows a similar strategy, special care is required to ensure that the Jacobian rows lie on appropriate tangent spaces. A hyperbolic LVM defines a stochastic mapping $f:\mathbb{H}^{D_{x}}_{\mathcal{L}}\rightarrow\mathbb{R}^{D_{y}}$, whose Jacobian is computed from the Riemannian gradients of $f$ as in (1). As for Riemannian submanifolds (Boumal, 2023), each Riemannian gradient $\mathrm{grad}_{\bm{x}}(f_{d})$ equals the orthogonal projection of the Euclidean gradient $\frac{\partial f_{d}}{\partial\bm{x}}\in\mathbb{R}^{D_{x}+1}$ onto $\mathcal{T}_{\bm{x}}\mathbb{H}^{D_{x}}_{\mathcal{L}}$,
@@
with $\bm{P}_{\bm{x}}\!=\!\bm{G}^{\mathcal{L}}+\bm{x}\bm{x}^{\mathsf{T}}$. Therefore, the Riemannian Jacobian $\tilde{\bm{J}}$ is computed from the Euclidean Jacobian $\bm{J}$ as ${\tilde{\bm{J}}\!=\![\bm{P}_{\bm{x}}\frac{\partial f_{1}}{\partial\bm{x}}\>\ldots\>\bm{P}_{\bm{x}}\frac{\partial f_{D_{y}}}{\partial\bm{x}}]^{\mathsf{T}}\!=\!\bm{J}\bm{P}^{\mathsf{T}}_{\bm{x}}}$. Assuming independent rows $\tilde{\bm{J}}_{d}$ as in the Euclidean case, the Jacobian distribution is computed from (8) as (Bishop, 2006, Chap. 2.3),
$p(\tilde{\bm{J}})=\prod_{d=1}^{D_{y}}\mathcal{N}(\tilde{\bm{J}}_{d}\mid\bm{P}_{\bm{x}}\bm{\mu}_{d},\bm{P}_{\bm{x}}\bm{\Sigma}_{J}\bm{P}_{\bm{x}}^{\mathsf{T}})\,.$ (12)
Similarly to $\bm{G}^{\text{P},\mathbb{R}}_{\bm{x}}$ in (9), the metric tensor $\bm{G}^{\text{P},\mathcal{L}}_{\bm{x}}$ follows a non-central Wishart distribution (Anderson, 1946),
@@
$\mathbb{E}[\bm{G}^{\text{P},\mathcal{L}}_{\bm{x}}]$ $=\mathbb{E}[\tilde{\bm{J}}]^{\mathsf{T}}\mathbb{E}[\tilde{\bm{J}}]+D_{y}\tilde{\bm{\Sigma}}_{J}$
$=\bm{P}_{\bm{x}}(\mathbb{E}[\bm{J}]^{\mathsf{T}}\mathbb{E}[\bm{J}]+D_{y}\bm{\Sigma}_{J})\bm{P}_{\bm{x}}^{\mathsf{T}}\,.$ (14)
Intuitively, the pullback metric $\bm{G}^{\text{P},\mathcal{L}}_{\bm{x}}$ is the orthogonal projection of the Euclidean pullback metric $\bm{G}^{\text{P},\mathbb{R}}_{\bm{x}}$ onto the tangent space $\mathcal{T}_{\bm{x}}\mathbb{H}^{D_{x}}_{\mathcal{L}}$ (see Fig. 1). It relates to the hyperbolic metric $\bm{G}^{\mathcal{L}}$ via the projection $\bm{P}_{\bm{x}}$, which is orthogonal with respect to the Lorentzian inner product.
@@ 3.2 The GPHLVM Pullback Metric
Here, we consider the case where the mapping $f$ is defined as a GPHLVM. We derive the distribution of the Riemannian Jacobian, which we then use to compute the expected GPHLVM pullback metric $\bm{G}^{\text{P},\mathcal{L}}_{\bm{x}^{*}}$. As for GPLVMs (Tosi et al., 2014), the joint distribution of the Jacobian $\tilde{\bm{J}}$ and observations $\bm{Y}=[\bm{y}_{1}\ldots\bm{y}_{N}]^{\mathsf{T}}\in\mathbb{R}^{N\times D_{y}}$ is given as,
$\mathcal{N}\left(\begin{bmatrix}\bm{Y}_{d}\\ \tilde{\bm{J}}_{d}\end{bmatrix}\Bigg{|}\begin{bmatrix}\bm{0}\\ \bm{0}\end{bmatrix},\begin{bmatrix}\bm{K}_{X}+\sigma_{y}^{2}\bm{I}_{N}&\partial k(\bm{X},\bm{x}^{*})\\ \partial k(\bm{x}^{*},\bm{X})&\partial^{2}k(\bm{x}^{*},\bm{x}^{*})\end{bmatrix}\right),$ (15)
@@
$\partial k(\bm{X},\bm{x}^{*})$ $=\frac{\partial}{\partial\bm{z}}k(\bm{X},\bm{z})\Big{|}_{\bm{z}=\bm{x}^{*}},$ (16)
$\partial^{2}k(\bm{x}^{*},\bm{x}^{*})$ $=\frac{\partial^{2}}{\partial\bm{z}\,\partial\bm{x}}k(\bm{x},\bm{z})\Big{|}_{\bm{x}=\bm{z}=\bm{x}^{*}}\,.$ (17)
While in the Euclidean case the kernel derivatives are straightforward, the hyperbolic setting is more challenging, as will be discussed in Sec. 4. Conditioning on the observations $\bm{Y}_{d}\in\mathbb{R}^{N}$ results in the probability distribution of the Jacobian at $\bm{x}^{*}$ (Bishop, 2006, Chap 2.3),
$p(\tilde{\bm{J}}\mid\bm{Y},\bm{X},\bm{x}_{*})$ $=\prod_{d=1}^{D_{y}}\mathcal{N}(\tilde{\bm{J}}_{d}\mid\bm{\mu}_{d},\bm{\Sigma}_{J}),$ (18)
$\text{with}\quad\quad\bm{\mu}_{d}$ $=\bm{S}_{J}\bm{Y}_{d}\,,$ (19)
$\bm{\Sigma}_{J}$ $=\partial^{2}k(\bm{x}^{*},\bm{x}^{*})-\bm{S}_{J}\partial k(\bm{X},\bm{x}^{*}),$ (20)
$\bm{S}_{J}$ $=\partial k(\bm{x}^{*},\bm{X})(\bm{K}_{X}+\sigma_{y}^{2}\bm{I}_{N})^{-1}\,.$ (21)
The expected pullback metric is computed using (14) as,
$\mathbb{E}[\bm{G}^{\text{P},\mathcal{L}}_{\bm{x}^{*}}]=\bm{P}_{\bm{x}^{*}}(\bm{\mu}_{J}^{\mathsf{T}}\bm{\mu}_{J}+D_{y}\bm{\Sigma}_{J})\bm{P}_{\bm{x}^{*}}^{\mathsf{T}}\,,$ (22)
with $\bm{\mu}_{J}=[\bm{\mu}_{1}\>\ldots\>\bm{\mu}_{D_{y}}]^{\mathsf{T}}\in\mathbb{R}^{D_{y}\times D_{x}}$.
@@ 3.3 Hyperbolic Pullback Geodesics
Having endowed the hyperbolic latent space with a pullback metric $\bm{G}^{\text{P},\mathcal{L}}$, we are now interested in computing shortest paths, i.e., geodesics, between latent hyperbolic points. While hyperbolic geodesics have a closed-form solution, they may not always align with the underlying data distribution (see Fig. 2). Instead, hyperbolic pullback geodesics naturally follow the data distribution, but require solving an optimization problem. Specifically, pullback geodesics are computed by minimizing the curve length, or equivalently the curve energy $E$ with respect to the pullback metric. When considering a discretized geodesic composed by a set of $M$ points $\bm{x}_{i}\in\mathbb{H}^{D_{x}}_{\mathcal{L}}$, this boils down to minimize,
@@
While it is possible to iteratively optimize the curve points $\bm{x}_{i}$ directly, this often leads to uneven spacing among them. This issue can be addressed either by using a parametric curve on the manifold (Gousenbourger et al., 2014) and optimize its parameters instead of the points directly, or by incorporating the spline energy as a regularization factor in the optimization process (Heeren et al., 2018). We follow the latter. The spline energy is $E_{\text{spline}}\!\approx\!\sum_{i=1}^{M-2}\text{d}_{\mathbb{H}^{D_{x}}_{\mathcal{L}}}(\bm{x}_{i},\bar{\bm{x}}_{i})^{2}$, where $\bar{\bm{x}}_{i}=\text{Exp}_{\bm{x}_{i-1}}\left(\frac{1}{2}\,\text{Log}_{\bm{x}_{i-1}}(\bm{x}_{i+1})\right)$ is the geodesic midpoint between $\bm{x}_{i-1}$ and $\bm{x}_{i+1}$. The final optimization problem is,
@@
with $\lambda$ weighting the influence of the regularization. As the optimization parameters are Riemannian, we leverage Riemannian optimizers such as Riemannian Adam (Bécigneul & Ganea, 2019) to optimize (23). The computation of pullback geodesics is summarised by Algorithm 1 in App. C.
Figure 2: Left and middle: Euclidean and hyperbolic pullback metrics on a $\mathsf{C}$-shape trajectory ( ). The geodesics of the base manifold ( ) and the pullbacks geodesic ( ) are depicted on both Euclidean and hyperbolic spaces. Right: Curve energy along the geodesics on the Euclidean (top) and hyperbolic (bottom) cases. These include the base manifold geodesic ( ), the base manifold geodesic with energy (23) evaluated using the pullback metric ( ), and the pullback geodesic ( ).
@@ 4 Hyperbolic Kernel Derivatives
In this section, we discuss the challenges encountered when using autodiff tools, in particular PyTorch  (Paszke et al., 2019), to compute the derivatives of the hyperbolic SE kernels. Table 1 shows that autodiff-based kernel derivatives are approximately five times slower than custom analytic implementations in the 2D case. More importantly, in the 3D case, autodiff fails to compute the derivatives entirely.
Table 1: Average computation times in seconds over $100$ runs, each evaluating the pullback metric $\bm{G}^{\text{P},\mathcal{L}}_{\bm{x}}$ and its derivatives at a random point $\bm{x}$. PyTorch fails to compute the 3D kernel derivatives.
@@
PyTorch derivatives $0.83\pm 0.06$-
Custom derivatives $\bm{0.16\pm 0.02}$ $\bm{0.06\pm 0.01}$
4.1 2D Hyperbolic SE Kernel Derivatives
@@
We compute the 2D hyperbolic SE kernel $k^{\mathbb{H}^{2}_{\mathcal{L}}}(\bm{x},\bm{z})$ via the Monte Carlo approximation (6). The pullback metric tensor $\bm{G}^{\text{P},\mathcal{L}}_{\bm{x}^{*}}$ is computed using the kernel derivatives $\frac{\partial}{\partial\bm{x}}k(\bm{x},\bm{z})|_{\bm{x}=\bm{x}^{*}}$ (16) and $\frac{\partial^{2}}{\partial\bm{z}\partial\bm{x}}k(\bm{x},\bm{z})|_{\bm{x}=\bm{z}=\bm{x}^{*}}$ (17). The only part of the kernel (6) that depends on the inputs $\bm{x},\bm{z}$ is the function $\Phi_{l}(\bm{x},\bm{z})$, whose derivatives are given by,
$\dfrac{\partial}{\partial\bm{x}}\Phi_{l}(\bm{x},\bm{z})$ $=\left[\dfrac{\partial}{\partial\bm{x}}\phi_{l}(\bm{x}_{\mathcal{P}})\right]\bar{\phi}_{l}(\bm{z}_{\mathcal{P}})\,,$ (25)
$\dfrac{\partial^{2}}{\partial\bm{z}\partial\bm{x}}\Phi_{l}(\bm{x},\bm{z})$ $=\left[\dfrac{\partial}{\partial\bm{x}}\phi_{l}(\bm{x}_{\mathcal{P}})\right]\left[\dfrac{\partial}{\partial\bm{z}}\bar{\phi}_{l}(\bm{z}_{\mathcal{P}})\right]^{\mathsf{T}}\,.$ (26)
The kernel derivatives (16)-(17) are thus straightforwardly determined from (25)-(26), and subsequently leveraged to compute $\bm{G}^{\text{P},\mathcal{L}}_{\bm{x}^{*}}$ using (22). Minimizing the curve energy (23) to compute pullback geodesics additionally requires the derivative $\frac{\partial}{\partial\bm{x}^{*}}\bm{G}^{\text{P},\mathcal{L}}_{\bm{x}^{*}}$. This, in turn, requires the derivative of the Jacobian of the covariance matrix $\frac{\partial}{\partial\bm{x}^{*}}\bm{\Sigma}_{J}$, which depends on the kernel derivatives $\frac{\partial^{3}}{\partial\bm{x}\partial\bm{z}\partial\bm{x}}k(\bm{x},\bm{z})|_{\bm{x}=\bm{z}=\bm{x}^{*}}$, and $\frac{\partial^{3}}{\partial\bm{z}^{2}\partial\bm{x}}k(\bm{x},\bm{z})|_{\bm{x}=\bm{z}=\bm{x}^{*}}$. Again, the kernel derivatives are effectively determined by the derivatives of $\Phi_{l}(\bm{x},\bm{z})$,
$\!\frac{\partial^{3}}{\partial\bm{x}\partial\bm{z}\partial\bm{x}}\Phi_{l}(\bm{x},\bm{z})\!$ $=\!\left[\!\dfrac{\partial^{2}}{\partial\bm{x}^{2}}\phi_{l}(\bm{x}_{\mathcal{P}}\!)\!\right]\!\times_{2}\!\left[\!\dfrac{\partial}{\partial\bm{z}}\bar{\phi}_{l}(\bm{z}_{\mathcal{P}}\!)\!\right],$ (27)
$\!\frac{\partial^{3}}{\partial\bm{z}^{2}\partial\bm{x}}\Phi_{l}(\bm{x},\bm{z})\!$ $=\!\left[\!\dfrac{\partial^{2}}{\partial\bm{z}^{2}}\bar{\phi}_{l}(\bm{z}_{\mathcal{P}}\!)\!\right]\!\times_{1}\!\left[\!\dfrac{\partial}{\partial\bm{x}}\phi_{l}(\bm{x}_{\mathcal{P}}\!)\!\right],$ (28)
where $\bm{A}\times_{i}\bm{B}$ refers to n-mode product (Kolda & Bader, 2009b) which multiplies the $i$-th dimension of $\bm{A}$ with the second dimension of $\bm{B}$. If $\bm{B}$ is a vector, $\bm{A}$ is unsqueezed at dimension $i$ before the multiplication. The matrices $\frac{\partial^{2}}{\partial\bm{x}^{2}}\phi_{l}(\bm{x}_{\mathcal{P}}),\frac{\partial^{2}}{\partial\bm{z}^{2}}\phi_{l}(\bm{z}_{\mathcal{P}})\in\mathbb{R}^{3\times 3}$ in (27)-(28) are therefore interpreted as $\mathbb{R}^{3\times 1\times 3}$ and $\mathbb{R}^{1\times 3\times 3}$, respectively. The complete derivation of all derivatives is provided in App. B.
4.2 3D Hyperbolic SE Kernel Derivatives
The 3D hyperbolic SE kernel is given by (5). In this section, we simplify its notation and write $k(\bm{x},\bm{z})=\frac{\rho}{s}e^{-\frac{\rho^{2}}{\nu}}$ by omitting the constant $\tau/C_{\infty}$, setting $\nu=2\kappa^{2}$, and introducing the helper variables $u=\langle\bm{x},\bm{z}\rangle_{\mathcal{L}}$ and $s=\sqrt{u^{2}-1}$. As in the 2D case, to obtain the expected pullback metric tensor (22), we compute the first two kernel derivatives,
$\dfrac{\partial}{\partial\bm{x}}\,k(\bm{x},\bm{z})$ $=g(u)\,\bm{G}^{\mathcal{L}}\bm{z}\,e^{-\frac{\rho^{2}}{\nu}}\,,$ (29)
$\dfrac{\partial^{2}}{\partial\bm{z}\partial{\bm{x}}}k(\bm{x},\bm{z})$ $=\left(h(u)\bm{G}^{\mathcal{L}}\bm{z}\bm{x}^{\mathsf{T}}\bm{G}^{\mathcal{L}}+g(u)\bm{G}^{\mathcal{L}}\right)e^{-\frac{\rho^{2}}{\nu}}\,,$ (30)

### v3

@@ 3 Metrics of Hyperbolic LVMs
Tosi et al. [51] introduced the pullback metric for Euclidean GPLVMs and its use to compute geodesics that adhere to the data distribution. Here, we extend this approach to hyperbolic LVMs, with GPHLVMs as a special case. Section 3.1 provides a general formulation of hyperbolic pullback metrics that applies to any hyperbolic LVM defined via an immersion $f:\mathbb{H}^{D_{x}}_{\mathcal{L}}\to\mathbb{R}^{D_{y}}$, provided that the Riemannian Jacobian of $f$ is available and follows a Gaussian distribution. Section 3.2 develops the hyperbolic pullback metric for a specific type of hyperbolic LVM, i.e., the GPHLVM, by deriving the probability distribution of the Riemannian Jacobian of the GPHLVM’s immersion function, which requires kernels and their derivatives. Finally, we compute hyperbolic pullback geodesics in Sec. 3.3.
@@ 3.1 A General Hyperbolic Pullback Metric
A deterministic immersion ${f:\mathbb{R}^{D_{x}}\to\mathbb{R}^{D_{y}}}$ pulls back the metric into the latent space following (5), i.e., ${\bm{G}^{\text{P},\mathbb{R}}_{\bm{x}}=\bm{J}^{\mathsf{T}}\bm{J}}$. In the context of LVMs, the immersion $f$ is stochastic. As developed in [51], the stochastic immersion $f$ induces a distribution over its Jacobian $\bm{J}$, which itself induces a distribution over the pullback metric. As [51], we consider LVMs for which the probability over $\bm{J}$ follows a Gaussian distribution. Assuming independent rows $\bm{J}_{d}\in\mathbb{R}^{D_{x}}$, each with its own mean but shared covariance matrix, the Jacobian distribution is of the form,
$p(\bm{J})=\prod_{d=1}^{D_{y}}\mathcal{N}(\bm{J}_{d}\mid\bm{\mu}_{\bm{J}_{d}},\bm{\Sigma}_{J})\,.$ (6)
Therefore, the metric tensor $\bm{G}^{\text{P},\mathbb{R}}_{\bm{x}}$ follows a non-central Wishart distribution [2],
@@
While the hyperbolic case follows a similar strategy, special care is required to ensure that the Jacobian rows lie on appropriate tangent spaces. Hyperbolic LVMs define a stochastic mapping $f:\mathbb{H}^{D_{x}}_{\mathcal{L}}\rightarrow\mathbb{R}^{D_{y}}$, whose Jacobian is formed by the Riemannian gradients of $f$ as in (1). As for Riemannian submanifolds [10], each Riemannian gradient $\mathrm{grad}_{\bm{x}}(f_{d})$ equals the orthogonal projection of the Euclidean gradient $\frac{\partial f_{d}}{\partial\bm{x}}\in\mathbb{R}^{D_{x}+1}$ onto $\mathcal{T}_{\bm{x}}\mathbb{H}^{D_{x}}_{\mathcal{L}}$,
@@
with $\bm{P}_{\bm{x}}\!=\!\bm{G}^{\mathcal{L}}+\bm{x}\bm{x}^{\mathsf{T}}$. Therefore, the Riemannian Jacobian $\tilde{\bm{J}}$ is computed from the Euclidean Jacobian $\bm{J}$ as ${\tilde{\bm{J}}\!=\![\bm{P}_{\bm{x}}\frac{\partial f_{1}}{\partial\bm{x}}\>\ldots\>\bm{P}_{\bm{x}}\frac{\partial f_{D_{y}}}{\partial\bm{x}}]^{\mathsf{T}}\!=\!\bm{J}\bm{P}^{\mathsf{T}}_{\bm{x}}}$. Assuming independent rows $\tilde{\bm{J}}_{d}$ as in the Euclidean case, the Jacobian distribution is computed from (6) as [6, Chap. 2.3],
$p(\tilde{\bm{J}})=\prod_{d=1}^{D_{y}}\mathcal{N}\big{(}\tilde{\bm{J}}_{d}\mid\bm{P}_{\bm{x}}\bm{\mu}_{\bm{J}_{d}},\bm{P}_{\bm{x}}\bm{\Sigma}_{J}\bm{P}_{\bm{x}}^{\mathsf{T}}\big{)}\,.$ (10)
Similarly to $\bm{G}^{\text{P},\mathbb{R}}_{\bm{x}}$ in (7), the metric tensor $\bm{G}^{\text{P},\mathcal{L}}_{\bm{x}}$ follows a non-central Wishart distribution [2],
@@
$\mathbb{E}[\bm{G}^{\text{P},\mathcal{L}}_{\bm{x}}]=\mathbb{E}[\tilde{\bm{J}}]^{\mathsf{T}}\mathbb{E}[\tilde{\bm{J}}]+D_{y}\tilde{\bm{\Sigma}}_{J}=\bm{P}_{\bm{x}}(\mathbb{E}[\bm{J}]^{\mathsf{T}}\mathbb{E}[\bm{J}]+D_{y}\bm{\Sigma}_{J})\bm{P}_{\bm{x}}^{\mathsf{T}}\,.$ (12)
Intuitively, the pullback metric $\bm{G}^{\text{P},\mathcal{L}}_{\bm{x}}$ is the orthogonal projection of the Euclidean pullback metric $\bm{G}^{\text{P},\mathbb{R}}_{\bm{x}}$ onto the tangent space $\mathcal{T}_{\bm{x}}\mathbb{H}^{D_{x}}_{\mathcal{L}}$ (see Fig. 1). It relates to the hyperbolic metric $\bm{G}^{\mathcal{L}}$ via the projection $\bm{P}_{\bm{x}}$, which is orthogonal with respect to the Lorentzian inner product. Note that the presented approach allows us to obtain pullback metrics for all hyperbolic LVMs where the Jacobian $\bm{J}$ of $f$ follows a Gaussian distribution, including GPHLVMs [25] and hyperbolic VAEs [38, 37].
@@ 3.2 The GPHLVM Pullback Metric
Here, we consider the case where the mapping $f$ is defined as a GPHLVM. We derive the distribution of the Riemannian Jacobian, which we then use to compute the expected GPHLVM pullback metric $\bm{G}^{\text{P},\mathcal{L}}_{\bm{x}^{*}}$. As for GPLVMs [51], the joint distribution of the Euclidean Jacobian $\bm{J}$ and observations $\bm{Y}=[\bm{y}_{1}\ldots\bm{y}_{N}]^{\mathsf{T}}\in\mathbb{R}^{N\times D_{y}}$ is given as,
$\mathcal{N}\left(\begin{bmatrix}\bm{Y}_{d}\\ \bm{J}_{d}\end{bmatrix}\Bigg{|}\begin{bmatrix}\bm{0}\\ \bm{0}\end{bmatrix},\begin{bmatrix}\bm{K}_{X}+\sigma_{y}^{2}\bm{I}_{N}&\partial k(\bm{X},\bm{x}^{*})\\ \partial k(\bm{x}^{*},\bm{X})&\partial^{2}k(\bm{x}^{*},\bm{x}^{*})\end{bmatrix}\right),$ (13)
@@
$\partial k(\bm{X},\bm{x}^{*})=\frac{\partial}{\partial\bm{z}}k(\bm{X},\bm{z})\Big{|}_{\bm{z}=\bm{x}^{*}}\quad\quad\text{and}\quad\quad\partial^{2}k(\bm{x}^{*},\bm{x}^{*})=\frac{\partial^{2}}{\partial\bm{z}\,\partial\bm{x}}k(\bm{x},\bm{z})\Big{|}_{\bm{x}=\bm{z}=\bm{x}^{*}}\,.$ (14)
While in the Euclidean case the kernel derivatives are straightforward, the hyperbolic setting is more challenging, as will be discussed in Sec. 4. Conditioning on the observations $\bm{Y}_{d}\in\mathbb{R}^{N}$ results in the probability distribution of the Jacobian at $\bm{x}^{*}$ [6, Chap 2.3],
$p(\bm{J}\mid\bm{Y},\bm{X},\bm{x}_{*})=\prod_{d=1}^{D_{y}}\mathcal{N}(\bm{J}_{d}\mid\bm{\mu}_{\bm{J}_{d}},\bm{\Sigma}_{J})\,\,\text{with}\,\,\bm{\mu}_{\bm{J}_{d}}=\bm{S}_{J}\bm{Y}_{d}\,\text{,}\,\bm{\Sigma}_{J}=\partial^{2}k(\bm{x}^{*},\bm{x}^{*})-\bm{S}_{J}\partial k(\bm{X},\bm{x}^{*}),$ (15)
and $\bm{S}_{J}=\partial k(\bm{x}^{*},\bm{X})(\bm{K}_{X}+\sigma_{y}^{2}\bm{I}_{N})^{-1}$. The distribution of the Riemannian Jacobian $\tilde{\bm{J}}$ is then obtained from (15) by using (10). The expected pullback metric is finally computed using (12) as,
$\mathbb{E}[\bm{G}^{\text{P},\mathcal{L}}_{\bm{x}^{*}}]=\bm{P}_{\bm{x}^{*}}(\bm{\mu}_{J}^{\mathsf{T}}\bm{\mu}_{J}+D_{y}\bm{\Sigma}_{J})\bm{P}_{\bm{x}^{*}}^{\mathsf{T}}\,,\quad\text{with}\quad\bm{\mu}_{J}=[\bm{\mu}_{\bm{J}_{1}}\>\ldots\>\bm{\mu}_{\bm{J}_{D_{y}}}]^{\mathsf{T}}\in\mathbb{R}^{D_{y}\times D_{x}}\,.$ (16)
@@ 3.3 Hyperbolic Pullback Geodesics
Vanilla geodesics on hyperbolic latent spaces naturally adhere to the hyperbolic geometry. However, they do not account for the underlying data structure and can traverse sparse data regions. To overcome this, we leverage pullback metrics on the hyperbolic manifold to design geodesics that comply with the geometry of both the hyperbolic space and the data distribution (see Figs. 1 and 2). Unlike hyperbolic geodesics that have a closed-form solution, hyperbolic pullback geodesics require solving an optimization problem. Specifically, pullback geodesics are computed by minimizing the curve length, or equivalently the curve energy $E$ with respect to the pullback metric. Considering a discretized geodesic composed by a set of $M$ points $\bm{x}_{i}\in\mathbb{H}^{D_{x}}_{\mathcal{L}}$, this boils down to minimize,
@@
While it is possible to iteratively optimize the curve points $\bm{x}_{i}$ directly, this often leads to uneven spacing among them. This issue can be addressed either by using a parametric curve on the manifold [20] and optimize its parameters instead of the points directly, or by incorporating the spline energy as a regularization factor in the optimization process [23]. We follow the latter. The spline energy is $E_{\text{spline}}\!\approx\!\sum_{i=1}^{M-2}\text{d}_{\mathbb{H}^{D_{x}}_{\mathcal{L}}}(\bm{x}_{i},\bar{\bm{x}}_{i})^{2}$, where $\bar{\bm{x}}_{i}=\text{Exp}_{\bm{x}_{i-1}}\left(\frac{1}{2}\,\text{Log}_{\bm{x}_{i-1}}(\bm{x}_{i+1})\right)$ is the geodesic midpoint between $\bm{x}_{i-1}$ and $\bm{x}_{i+1}$. The final optimization problem is,
@@
with $\lambda$ weighting the influence of the regularization. As the optimization parameters are Riemannian, we leverage Riemannian optimizers such as Riemannian Adam [11] to optimize (17). The computation of pullback geodesics is summarized in Algorithm 1 in App. C.
Figure 2: Left and middle: Euclidean and hyperbolic pullback metrics on an embedded $\mathsf{C}$-shape trajectory ( ) with base manifold ( ) and pullback ( ) geodesics. Right: Curve energy along the geodesics on the Euclidean (top) and hyperbolic (bottom) cases. These include the base manifold geodesic ( ), the base manifold geodesic with energy (17) evaluated using the pullback metric ( ), and the pullback geodesic ( ).
@@ 4 Hyperbolic Kernel Derivatives
Here we discuss the challenges of computing derivatives for hyperbolic SE kernels for which standard autodiff tools cannot be used directly. Table 1 shows that autodiff-based kernel derivatives computed with PyTorch [42] are approximately five times slower than analytic implementations in the 2D case. In the 3D case, autodiff fails to compute the derivatives entirely. To overcome these issues, we derive analytic kernel derivatives and focus our analysis on the 2D and 3D cases. This choice is motivated
Table 1: Average computation times in seconds over $100$ runs, each evaluating the pullback metric $\bm{G}^{\text{P},\mathcal{L}}_{\bm{x}}$ and its derivatives at a random point $\bm{x}$.
@@
PyTorch $0.83\pm 0.06$-
Analytic $\bm{0.16\pm 0.02}$ $\bm{0.06\pm 0.01}$
by the increased capacity of the hyperbolic manifold to embed hierarchical data in low-dimensional spaces, coupled with the computational efficiency and ease of visualization of 2D and 3D latent spaces. Note that the forthcoming derivations could serve as a basis for higher-dimensional hyperbolic kernels defined as derivatives of 2D and 3D kernels (3[21].
@@
2D Hyperbolic SE Kernel Derivatives: We compute the 2D hyperbolic SE kernel $k^{\mathbb{H}^{2}_{\mathcal{L}}}(\bm{x},\bm{z})$ via the Monte Carlo approximation (4). The pullback metric tensor $\bm{G}^{\text{P},\mathcal{L}}_{\bm{x}^{*}}$ is computed in (16) using the kernel derivatives $\frac{\partial k(\bm{x},\bm{z})}{\partial\bm{x}}|_{\bm{x}=\bm{x}^{*}}$ and $\frac{\partial^{2}k(\bm{x},\bm{z})}{\partial\bm{z}\partial\bm{x}}|_{\bm{x}=\bm{z}=\bm{x}^{*}}$ (14). The only part of the kernel (4) that depends on the inputs $\bm{x},\bm{z}$ is the function $\Phi_{l}(\bm{x},\bm{z})$, whose derivatives are given in App. B.1. Minimizing the curve energy (17) to compute pullback geodesics additionally requires the derivative $\frac{\partial}{\partial\bm{x}^{*}}\bm{G}^{\text{P},\mathcal{L}}_{\bm{x}^{*}}$. This, in turn, requires the derivative of the Jacobian of the covariance matrix $\frac{\partial\bm{\Sigma}_{J}}{\partial\bm{x}^{*}}$, which depends on the kernel derivatives $\frac{\partial^{3}k(\bm{x},\bm{z})}{\partial\bm{x}\partial\bm{z}\partial\bm{x}}|_{\bm{x}=\bm{z}=\bm{x}^{*}}$, and $\frac{\partial^{3}k(\bm{x},\bm{z})}{\partial\bm{z}^{2}\partial\bm{x}}|_{\bm{x}=\bm{z}=\bm{x}^{*}}$. Again, the kernel derivatives are determined by the derivatives of $\Phi_{l}(\bm{x},\bm{z})$. The complete derivation of all derivatives is provided in App. B.1.